OpenStudy (anonymous):

evaluate if a=2 and b=-5 ab^3

OpenStudy (anonymous):

-250

OpenStudy (angela210793):

all you have to do is put 2 instead of a and -5 instead of b---> 2*(-5)^3=2*(-125)=-250